# Binary superfluids: low-energy properties and dissipative processes from spontaneous emission of massive phonons

### v1

@@ III.1 Lagrangian of binary bosonic systems
as each component has distinct chemical potential and velocity. Here, $\gamma_{A/B}=(1-v_{A/B}^{2})^{-1/2}$ are the Lorentz factors of each fluid component. We consider only quartic intra- and inter-species interactions, $V_{U}$, to which a Rabi term, $V_{\lambda}$, is conveniently added:
@@
where $b_{l,i}$ are the ladder operators, the indices $(i,j)$ label the two bosonic flavors, that in principle can have different numbers or fillings, while the indices $(l,m)$ refers to the position in a lattice or in continuous space. The added one-particle potential, $V_{\text{trap}}(\mathbf{r}_{l})$, is typically harmonic and is necessary to provide particle trapping.
@@
Referring back to Eq. (III.1), the $U_{ij}$ quartic couplings determine the strengths of the density-density boson interactions. Perturbation theory holds when $|U_{ij}|\ll 1$, moreover for
@@
while the sign of the inter-flavor coupling, $U_{AB}$, is not fixed.
@@ III.2 Symmetries
with ${\cal I}_{2}$ the $2\times 2$ identity, and $A_{\nu}$ the hermitian $2\times 2$ matrix that represents a general external vector field. As in Eqs. (30) and (31), we assume that it has only time-like components that are diagonal in the flavor basis. Indicating with $\sigma_{i}$ the Pauli matrices, the only nonvanishing terms in $A_{\nu}$ can be cast as
@@
where $SU(2)$ generates flavor rotations. The above relation simply states that for two fluids having the same chemical potentials and velocities, the kinetic term does not distinguish the two fluids, thus it is invariant under a global phase redefinition and flavor rotations.
In the vector notation, the mass –or quadratic– term contains also the Rabi interaction and reads
@@
  • if $\lambda=0$ and $\Delta m^{2}=0$, the mass term symmetry is $G_{T}$: the two fields are indistinguishable and invariant under global phase and flavor rotations;
@@
  • if $\lambda=0$ and $\Delta m^{2}\neq 0$, the mass term symmetry is $G$: the two fields are distinguishable because they have different masses;
@@
In this work, we will focus on the last two cases. For sufficiently small values of $\lambda$, the perturbative Rabi term produces a soft breaking of the axial symmetry. As a consequence, in the spontaneously broken phase, we expect to have one true NGB, originating from the breaking of the $U(1)_{\cal V}$ symmetry, plus a pseudo-NBG associated to the breaking of the non-exact $U(1)_{\cal A}$ symmetry: a massless phonon and a massive phonon. In the remaining of this paper, we will consider the corrections induced by the Rabi term only at leading order ${\cal O}(\lambda)$, keeping in mind that one-loop corrections are of order $\lambda\log(\lambda)$Li and Pagels (1971).
We consider now the density-density interactions in Eq. (III.1): in the vector basis both the inter-species and intra-species interactions can be rewritten as
@@
the two flavor projectors. Since the interaction terms are proportional to ${\cal I}_{2}$ and $\sigma_{3}$, they are invariant under $G$ but not under the enlarged group $G_{\text{T}}$: they distinguish the two bosonic species. The terms proportional to $\sigma_{3}$ vanish only when $U_{AA}$ = $U_{BB}$, and $U_{AA}=2|U_{AB}|$, however, this condition violates the stability requirement in Eq. (36).
Let us briefly recap the most important aspects of the above analysis. For vanishing Rabi coupling, the Lagrangian has the $G$ symmetry given in Eq. (42). If this symmetry is spontaneously-broken the system is expected to have two genuine NGBs, that is two massless phonons. If $\lambda\neq 0$, the system has only the $U(1)_{\cal V}$ symmetry and in the superfluid ground-state only one NGB can appear. Furthermore, if the Rabi term can be treated as a soft explicit symmetry breaking, the low-energy theory is expected to include a pseudo-NGB, with a square mass proportional to $\lambda$.
@@ IV Pressure of binary superfluids
The considered system consists of two interacting gases having both intra-species and inter-species interactions, as well as a perturbative Rabi coupling. In absence of both the inter-species and Rabi interactions, the two fluids would be non-interacting and the total pressure of the system would be equal to the sum of two partial ones associated to each flavor. By adding the interactions between the two components, we might naively expect that the total pressure be cannot be separated into two independent contributions. Nevertheless, if the Rabi coupling vanishes and the $G$ symmetry holds, the ground-state Lagrangian can be cast as a sum of two terms, and the total pressure is the sum of two partial pressures. Furthermore, the same separation can be extended for perturbative Rabi coupling. In the following, we investigate which are the conditions for such separation.
@@ IV.1 Flavor rotation in the ground state
the Lagrangian in Eq. (29) with the potential in Eq. (III.1) reads
@@
where $v^{\mu}_{A}$ and $v^{\mu}_{B}$ are the two fluids four-velocities. Hence, from Eq. (IV.1), we obtain the background Lagrangian
@@
Before estimating the mean field values of the radial fields, we remark a subtle aspect about the phase fields. They are not variational parameters, as they are connected by Eq. (53) with the fluids’ flow (the same holds in the single-fluid case). Since the Rabi coupling explicitly depends on the relative phase, it tends to fix such a difference. For instance, in the broken phase, the mean field values $\rho_{A}$ and $\rho_{B}$ are positive, thus if $\lambda<0$ the background pressure is maximized for
@@ IV.2 Vanishing Rabi coupling
and are the analogous of the expression for the single scalar field in Eq. (13). In the $I,\,II$ basis the conditions to have nonvanishing expectation values of the scalar fields $\rho_{I}$ and $\rho_{II}$, are $\mu_{I}>m_{I}$ and $\mu_{II}>m_{II}$, respectively. When expressing these conditions in the $A,B$ basis, we have that
@@
$\tan\alpha>-\frac{\mu_{A}^{2}-m_{A}^{2}}{\mu_{B}^{2}-m_{B}^{2}}\quad\text{{\color[rgb]{0,0,0}and}}\quad\tan\alpha>\frac{\mu_{B}^{2}-m_{B}^{2}}{\mu_{A}^{2}-m_{A}^{2}}\,,$ (80)
@@
it is possible that $\rho_{I}>0$, while $\rho_{II}=0$. In the following, we will assume for simplicity that the second condition in (80) is satisfied, therefore both $\rho_{I}$ and $\rho_{II}$ are positive.
@@ IV.3 Including the Rabi coupling
and is a non-trivial coupling between the $\rho_{I}$ and $\rho_{II}$ fields. If the explicit breaking is small, we can treat $V_{\lambda}$ as a perturbation. In this case, we expand the mean field solutions at leading order in $\lambda$ as follows
@@
$\frac{|\lambda_{AB}|}{\rho_{I}^{2}}\ll 1\quad\text{and}\quad\frac{|\lambda_{AB}|}{\rho_{II}^{2}}\ll 1\,.$ (87)
@@
$\delta\rho_{I}=-\frac{\rho_{I}^{3}\sin(2\alpha)-\rho_{I}\rho_{II}^{2}}{U_{+}\sqrt{2(\rho_{I}^{4}-\rho_{II}^{4})\sin(2\alpha)-4\rho_{I}^{2}\rho_{II}^{2}\cos(2\alpha)}}\,,$ (88)
@@
$\delta\rho_{II}=+\frac{\rho_{II}^{3}\sin(2\alpha)+\rho_{II}\rho_{I}^{2}}{U_{-}\sqrt{2(\rho_{I}^{4}-\rho_{II}^{4})\sin(2\alpha)-4\rho_{I}^{2}\rho_{II}^{2}\cos(2\alpha)}}\,.$ (89)
@@
This, in turn, implies that $\rho_{A}\simeq\rho_{B}$ (see Eq. (73)) and thus
@@
Comparing these expressions with the estimate in Eq. (35), we better comprehend the condition to treat perturbatively the Rabi term. In the broken phase, the actual dimensionless expansion parameter is the ratio between the Rabi coupling and the expectation values of the rotated radial fields. Since $\rho_{I,II}$ are of the order of $m_{A,B}$, the condition $\tilde{\lambda}\ll\tilde{U}_{ij}$ holds.
@@
We remark that for both non-vanishing $U_{AB}$ and $\lambda$, in general, it is not possible to simultaneously diagonalize $V_{U}$ and $V_{\lambda}$. They differ in the dependence in $\rho_{A}$ and $\rho_{B}$: quadratic for the inter-species interaction, and linear for the Rabi term. Consequently, the diagonalization of $V_{U}$ requires a non linear transformation of the fields, as in Eq. (IV.1), leaving the Rabi term non diagonal. The specular situation occurs when the Rabi term dominates over the inter-species interaction. If we diagonalize the Rabi term, the transformed fields remain interacting due to a non-vanishing $U_{AB}$epori et al. (2018).
@@ V Quantum fluctuations in binary superfluids
In the previous section, we characterized the mean field Lagrangian of the boson mixture. Following the same steps of the single-fluid model, we analyze the effective action for the longwavelength excitations. We replace
@@
and throughout the remaining on this work, $\rho_{i}$ and $\theta_{i}$ indicate the mean field solutions, while $\tilde{\rho}_{i}$ and $\tilde{\theta}_{i}$ are their corresponding fluctuations. We expand the Lagrangian as
@@
with ${\cal L}_{0}$ the mean-field one in Eq. (54), while ${\cal L}_{2}$ is quadratic in the fields’ fluctuations. In the following sections, we will discuss the case for vanishing Rabi interaction ($\lambda=0$), and the non-vanishing case ($\lambda\neq 0$). We will show that $\mathcal{L}_{2}$ can be written as the one of two non-interacting scalar fields, propagating on an emergent acoustic metric. The Rabi term will induce an effective mass in one of the scalar fields.
@@ V.1 Fluctuations with vanishing Rabi term
We diagonalize $\mathcal{L}_{2}$ with a suitable transformation in flavor space of the fields $\hat{\bm{\rho}}$ and $\hat{\bm{\theta}}$. The first and last term in the r.h.s. of Eq. (108) are separately invariant under any flavor mixing transformation of $\hat{\bm{\rho}}$ and $\hat{\bm{\theta}}$, as the one in Eq. (63). However, the invariance of the second term locks the mixing angle to be the same for $\hat{\bm{\rho}}$ and $\hat{\bm{\theta}}$. The locking implies that the vectors in the new basis are
@@
are not the fluctuations of $\rho_{I/II}$ and $\theta_{I/II}$. They can be written as a linear combination of $\tilde{\rho}_{a/b}$ and $\tilde{\theta}_{a/b}$, respectively.
In the ${a,b}$ basis the matrix $H$ is diagonal, thus no mixing exists, in other words
@@
and ${\cal L}_{2b}$ is given by an analogous expression. They have the same formal expression obtained in the single-fluid case, see Eq. (19), when neglecting the kinetic term of the radial mode. Note that the radial modes have dimension $2$, therefore their masses
@@
and given the expression of the $H$–matrix components in (V.1), each sound mode depends on the intra- and inter-species couplings, as well as on the chemical potentials $\mu_{A}$ and $\mu_{B}$ and the masses $m_{A}$ and $m_{B}$.
@@ V.2 Including the Rabi term in the fluctuations
and it produces a pletora of terms. The linear terms in the fluctuations vanish at the stationary point. There are two quadratic terms in the fluctuations, which have distinct effects. One term is proportional to $\tilde{\rho}_{A}\tilde{\rho}_{B}$ and thus contributes to the $H_{12}$ matrix element, see Eq. (V.1). This produces a trivial small shift of the effective masses of the radial fluctuations. The second quadratic term in the fluctuations is
@@
Before doing that, we observe that to have the total pressure as the sum of two partial pressure in the background analysis, we had to assume that $\rho_{I}\gg\rho_{II}$ and, correspondingly, a background flavor rotation angle close to maximal mixing, that is $\alpha\simeq\pi/4$. We expect that the same conditions would lead to an effective theory with two decoupled modes, one massless and one massive. To this end, we follow the same procedure discussed for vanishing Rabi coupling. In particular, we rescale the phases as in Eq. (102) and obtain that, close to maximal background mixing, the condition $\rho_{I}\gg\rho_{II}$ implies $\rho_{B}\simeq\rho_{A}$, see Eq. (72). Then, Eq. (119) can be rewritten as
@@
Upon integrating out the radial fluctuations using the same procedure discussed for the single-fluid case in Sec. II.2, we obtain the quadratic effective Lagrangian in the phases fluctuations
@@
the mass squared of the $\tilde{\theta}_{b}$ phonon. A similar expression was obtained in Liberati et al. (2006a) with a different method. The above expression implies that $\lambda\cos(\theta_{A}-\theta_{B})$ must be negative, otherwise the $\tilde{\theta}_{b}$ field would be tachyonic. This is the same condition needed to have a positive contribution of the Rabi term to the background pressure, see Eq. (54). Taking $\lambda<0$, the Rabi pressure is maximal for $\theta_{A}=\theta_{B}$, see Eq. (55), which is also the condition that maximizes the mass of the $\tilde{\theta}_{b}$ field. Alternatively, the favored state may be a crystal of sine-Gordon kinks, see Eq. (52), and the negative value of the squared mass would actually indicate the presence of phonon excitations of the crystal Takayama and Oka (1993).
An interesting point is that, from Eq. (53), the difference of the background phases is related to a difference in the chemical potentials and fluid velocities. Since in the derivation we have assumed that the two velocities are equal (it was not a necessary condition but it helped to simplify the analysis), the condition to have a positive defined squared mass can be related to a condition on the chemical potentials. As we have previously discussed, it is possible to prepare the system in such a way that
@@
where $\Theta_{0}$ is some constant phase that can be reabsorbed in the definition of the background phases. Then, the Rabi coupling will tend to make the two chemical potentials equal; indeed any chemical potential difference would produce a mode that oscillates in time between a stable and a tachyonic configuration, signaling that the system is not in the ground-state.

### v2

@@ III.1 Lagrangian of binary boson systems
as each component has distinct chemical potential and velocity. Here, $\gamma_{A/B}=(1-v_{A/B}^{2})^{-1/2}$ are the Lorentz factors of each fluid component. We consider only quartic intraspecies and interspecies interactions, $V_{U}$, to which a Rabi term, $V_{\lambda}$, is conveniently added:
@@
where $b_{l,i}$ are the ladder operators, the indices $(i,j)$ label the two boson flavors, that in principle can have different numbers or fillings, while the indices $(l,m)$ refers to the position in a lattice or in continuous space. The added one-particle potential, $V_{\text{trap}}(\mathbf{r}_{l})$, is typically harmonic and is necessary to provide particle trapping.
@@
Referring back to Eq. (III.1), the $U_{ij}$ quartic couplings determine the strengths of the density-density boson interactions. Perturbation theory holds when $|U_{ij}|\ll 1$; moreover for
@@
while the sign of the interflavor coupling, $U_{AB}$, is not fixed.
@@ III.2 Symmetries
with ${\cal I}_{2}$ the $2\times 2$ identity, and $A_{\nu}$ the Hermitian $2\times 2$ matrix that represents a general external vector field. As in Eqs. (32) and (33), we assume that it has only timelike components that are diagonal in the flavor basis. Indicating with $\sigma_{i}$ the Pauli matrices, the only nonvanishing terms in $A_{\nu}$ can be cast as
@@
where $SU(2)$ generates flavor rotations. The above relation simply states that for two fluids having the same chemical potentials and velocities, the kinetic term does not distinguish the two fluids; thus it is invariant under a global phase redefinition and flavor rotations.
In the vector notation, the mass–or quadratic–term contains also the Rabi interaction and reads as
@@
  1. 1.
    If $\lambda=0$ and $\Delta m^{2}=0$, the mass term symmetry is $G_{T}$: the two fields are indistinguishable and invariant under global phase and flavor rotations.
  2. 2.
@@
  3. 3.
    If $\lambda=0$ and $\Delta m^{2}\neq 0$, the mass term symmetry is $G$: the two fields are distinguishable because they have different masses.
  4. 4.
@@
In this work, we will focus on the last two cases. For sufficiently small values of $\lambda$, the perturbative Rabi term produces a soft breaking of the axial symmetry. As a consequence, in the spontaneously broken phase, we expect to have one true NGB, originating from the breaking of the $U(1)_{\cal V}$ symmetry, plus a pseudo-NBG associated to the breaking of the nonexact $U(1)_{\cal A}$ symmetry: a massless phonon and a massive phonon. In the remainder of this paper, we will consider the corrections induced by the Rabi term only at order ${\cal O}(\lambda)$, ignoring the possible formation of topological solitons; see however Li and Pagels (1971).
We discuss now the density-density interactions in Eq. (III.1): in the vector basis both the interspecies and intraspecies interactions can be rewritten as
@@
the two flavor projectors. Since the interaction terms are proportional to ${\cal I}_{2}$ and $\sigma_{3}$, they are invariant under $G$ but not under the enlarged group $G_{\text{T}}$: they distinguish the two boson species. The terms proportional to $\sigma_{3}$ vanish only when $U_{AA}$ = $U_{BB}$, and $U_{AA}=2|U_{AB}|$, however, this condition violates the stability requirement in Eq. (38).
Let us briefly recap the most important aspects of the above analysis. For vanishing Rabi coupling, the Lagrangian has the $G$ symmetry given in Eq. (44). If this symmetry is spontaneously broken the system is expected to have two genuine NGBs, that is two massless phonons. If $\lambda\neq 0$, the system has only the $U(1)_{\cal V}$ symmetry and in the superfluid ground state only one NGB can appear. Furthermore, if the Rabi term can be treated as a soft explicit symmetry breaking, the low-energy theory is expected to include a pseudo-NGB, with a square mass proportional to $\lambda$.
@@ IV Pressure of binary superfluids
The considered system consists of two interacting gases having both intraspecies and interspecies interactions, as well as a perturbative Rabi coupling. In absence of both the interspecies and Rabi interactions, the two fluids would be noninteracting and the total pressure of the system would be equal to the sum of two partial ones associated to each flavor. By adding the interactions between the two components, we might naively expect that the total pressure cannot be separated into two independent contributions. Nevertheless, if the Rabi coupling vanishes and the $G$ symmetry holds, the ground-state Lagrangian can be cast as a sum of two terms, and the total pressure is the sum of two partial pressures. Furthermore, the same separation can be extended for perturbative Rabi coupling. In the following, we investigate which are the conditions for such separation.
@@ IV.1 Flavor rotation in the ground state
the Lagrangian in Eq. (31) with the potential in Eq. (III.1) reads as
@@
where $v^{\mu}_{A}$ and $v^{\mu}_{B}$ are the two fluid four-velocities. Hence, from Eq. (IV.1), we obtain the background Lagrangian
@@
Before estimating the mean field values of the radial fields, we remark on a subtle aspect about the phase fields. They are not variational parameters, as they are connected by Eq. (55) with the fluids’ flow (the same holds in the single-fluid case). Since the Rabi coupling explicitly depends on the relative phase, it tends to fix such a difference. For instance, in the broken phase, the mean field values $\rho_{A}$ and $\rho_{B}$ are positive; thus if $\lambda<0$ the background pressure is maximized for
@@ IV.2 Vanishing Rabi coupling
and are analogous to the expression for the single scalar field in Eq. (13). In the $I,\,II$ basis the conditions to have nonvanishing expectation values of the scalar fields $\rho_{I}$ and $\rho_{II}$, are $\mu_{I}>m_{I}$ and $\mu_{II}>m_{II}$, respectively. When expressing these conditions in the $A,B$ basis, we have that
@@
$\tan\alpha>-\frac{\mu_{A}^{2}-m_{A}^{2}}{\mu_{B}^{2}-m_{B}^{2}}\quad\text{and}\quad\tan\alpha>\frac{\mu_{B}^{2}-m_{B}^{2}}{\mu_{A}^{2}-m_{A}^{2}}\,,$ (82)
@@
it is possible that $\rho_{I}>0$, while $\rho_{II}=0$. In the following, we will assume for simplicity that the second condition in (82) is satisfied; therefore both $\rho_{I}$ and $\rho_{II}$ are positive.
@@ IV.3 Including the Rabi coupling
and it is a nontrivial coupling between the $\rho_{I}$ and $\rho_{II}$ fields. If the explicit breaking is small, we can treat $V_{\lambda}$ as a perturbation. In this case, we expand the mean field solutions at leading order in $\lambda$ as follows
@@
$\frac{|\lambda_{AB}|}{\rho_{I}^{2}}\ll 1\quad\text{and}\quad\frac{|\lambda_{AB}|}{\rho_{II}^{2}}\ll 1\,,$ (89)
@@
$\delta\rho_{I}=-\frac{\rho_{I}^{3}\sin(2\alpha)-\rho_{I}\rho_{II}^{2}\cos(2\alpha)}{U_{+}\sqrt{2(\rho_{I}^{4}-\rho_{II}^{4})\sin(2\alpha)-4\rho_{I}^{2}\rho_{II}^{2}\cos(2\alpha)}}\,,$ (90)
@@
$\delta\rho_{II}=+\frac{\rho_{II}^{3}\sin(2\alpha)+\rho_{II}\rho_{I}^{2}\cos(2\alpha)}{U_{-}\sqrt{2(\rho_{I}^{4}-\rho_{II}^{4})\sin(2\alpha)-4\rho_{I}^{2}\rho_{II}^{2}\cos(2\alpha)}}\,.$ (91)
@@
This, in turn, implies that $\rho_{A}\simeq\rho_{B}$ [see Eq. (75)] and thus
@@
Comparing these expressions with the estimate in Eq. (37), we better comprehend the condition for treating perturbatively the Rabi term. In the broken phase, the actual dimensionless expansion parameter is the ratio between the Rabi coupling and the expectation values of the rotated radial fields. Since $\rho_{I,II}$ are of the order of $m_{A,B}$, the condition $\tilde{\lambda}\ll\tilde{U}_{ij}$ holds.
@@
We remark that for both nonvanishing $U_{AB}$ and $\lambda$, in general, it is not possible to simultaneously diagonalize $V_{U}$ and $V_{\lambda}$. They differ in the dependence in $\rho_{A}$ and $\rho_{B}$: quadratic for the interspecies interaction, and linear for the Rabi term. Consequently, the diagonalization of $V_{U}$ requires a non linear transformation of the fields, as in Eq. (IV.1), leaving the Rabi term non diagonal. The specular situation occurs when the Rabi term dominates over the interspecies interaction. If we diagonalize the Rabi term, the transformed fields remain interacting due to a nonvanishing $U_{AB}$epori et al. (2018).
@@ V Quantum fluctuations in binary superfluids
In the previous section, we characterized the mean field Lagrangian of the boson mixture. Following the same steps of the single-fluid model, we analyze the effective action for the long-wavelength excitations. We replace
@@
and throughout the remainder of this work, $\rho_{i}$ and $\theta_{i}$ indicate the mean field solutions, while $\tilde{\rho}_{i}$ and $\tilde{\theta}_{i}$ are their corresponding fluctuations. We expand the Lagrangian as
@@
with ${\cal L}_{0}$ the mean-field one in Eq. (56), while ${\cal L}_{2}$ is quadratic in the fields’ fluctuations. In the following sections, we first discuss the case for vanishing Rabi interaction ($\lambda=0$), and then the nonvanishing case ($\lambda\neq 0$). We will show that $\mathcal{L}_{2}$ can be written as the one of two noninteracting scalar fields, propagating on an emergent acoustic metric. The Rabi term will induce an effective mass in one of the scalar fields.
@@ V.1 Fluctuations with vanishing Rabi term
We diagonalize $\mathcal{L}_{2}$ with a suitable transformation in flavor space of the fields $\hat{\bm{\rho}}$ and $\hat{\bm{\theta}}$. The first and last term in the rhs of Eq. (110) are separately invariant under any flavor mixing transformation of $\hat{\bm{\rho}}$ and $\hat{\bm{\theta}}$, as the one in Eq. (65). However, the invariance of the second term locks the mixing angle to be the same for $\hat{\bm{\rho}}$ and $\hat{\bm{\theta}}$. The locking implies that the vectors in the new basis are
@@
are not the fluctuations of $\rho_{I/II}$ and $\theta_{I/II}$. However, they can be written as a linear combination of $\tilde{\rho}_{a/b}$ and $\tilde{\theta}_{a/b}$, respectively.
In the ${a,b}$ basis the matrix $H$ is diagonal; thus no mixing exists. In other words
@@
and ${\cal L}_{2b}$ is given by an analogous expression. They have the same formal expression obtained in the single-fluid case [see Eq. (19)] when neglecting the kinetic term of the radial mode. Note that the radial modes have dimension $2$; therefore their masses
@@
and given the expression of the $H$ matrix components in (V.1), each sound mode depends on the intraspecies and interspecies couplings, as well as on the chemical potentials $\mu_{A}$ and $\mu_{B}$ and the masses $m_{A}$ and $m_{B}$.
@@ V.2 Including the Rabi term in the fluctuations
and it produces a plethora of terms. The linear terms in the fluctuations vanish at the stationary point. There are two quadratic terms in the fluctuations, which have distinct effects. One term is proportional to $\tilde{\rho}_{A}\tilde{\rho}_{B}$ and thus contributes to the $H_{12}$ matrix element, see Eq. (V.1). This produces a trivial small shift of the effective masses of the radial fluctuations. The second quadratic term in the fluctuations is
@@
Before doing that, we observe that to have the total pressure as the sum of two partial pressures in the background analysis, we had to assume that $\rho_{I}\gg\rho_{II}$ and, correspondingly, a background flavor rotation angle close to maximal mixing, that is $\alpha\simeq\pi/4$. We expect that the same conditions would lead to an effective theory with two decoupled modes, one massless and one massive. To this end, we follow the same procedure discussed for vanishing Rabi coupling. In particular, we rescale the phases as in Eq. (104) and obtain that, close to maximal background mixing, the condition $\rho_{I}\gg\rho_{II}$ implies $\rho_{B}\simeq\rho_{A}$, see Eq. (74). Then, Eq. (121) can be rewritten as
@@
Upon integrating out the radial fluctuations using the same procedure discussed for the single-fluid case in Sec. II.2, we obtain the quadratic effective Lagrangian in the phase fluctuations
@@
the mass squared of the $\tilde{\theta}_{b}$ phonon. A similar expression was obtained in Liberati et al. (2006a) with a different method. The above expression implies that $\lambda\cos(\theta_{A}-\theta_{B})$ must be negative; otherwise the $\tilde{\theta}_{b}$ field would be tachyonic. This is the same condition needed to have a positive contribution of the Rabi term to the background pressure, see Eq. (56). Taking $\lambda<0$, the Rabi pressure is maximal for $\theta_{A}=\theta_{B}$ [see Eq. (57)], which is also the condition that maximizes the mass of the $\tilde{\theta}_{b}$ field. Alternatively, the favored state may be a crystal of sine-Gordon kinks [see Eq. (54)], and the negative value of the squared mass would actually indicate the presence of phonon excitations of the crystal Takayama and Oka (1993).
An interesting point is that, from Eq. (55), the difference of the background phases is related to a difference in the chemical potentials and fluid velocities. Since in the derivation we have assumed that the two velocities are equal (it was not a necessary condition, but it helped to simplify the analysis), the condition to have a positive defined squared mass can be related to a condition on the chemical potentials. As we have previously discussed, it is possible to prepare the system in such a way that
@@
where $\Theta_{0}$ is some constant phase that can be reabsorbed in the definition of the background phases. Then, the Rabi coupling will tend to make the two chemical potentials equal; indeed any chemical potential difference would produce a mode that oscillates in time between a stable and a tachyonic configuration, signaling that the system is not in the ground state.